# Hierarchical certification of non-classical network correlations

### v3

@@ Abstract
With the increased availability of quantum technological devices, it becomes more important to have tools to guarantee their correct non-classical behavior. This is especially important for quantum networks, which constitute the platforms where multipartite cryptographic protocols will be implemented, and where guarantees of non-classicality translate into security proofs. We derive linear and non-linear Bell-like inequalities for networks, whose violation certifies the absence of a minimum number of classical sources in them. We do so, firstly, without assuming that nature is ultimately governed by quantum mechanics, providing a hierarchy interpolating between network nonlocality and full network nonlocality. Secondly we insert this assumption, which leads to results more amenable to certification in experiments.
@@
Figure 1: The main networks discussed in this work. (a) is the network underlying simple entanglement swapping experiments. Correlations generated in the network are considered NN if they cannot be generated as in the top-left setup, FNN if they cannot be generated as in the bottom-left setup, and FQNN if they cannot be generated as in the top-right setup. In the appendices we show that realizations in the bottom-right setup produce all these types of correlations. (b) is a generalization of the previous scenario to chains, and (c) is its generalization to stars.
Nonlinear Bell inequalities have been used to study network nonlocality (NN), especially in the tripartite network underlying entanglement swapping depicted in Fig. 1(a) [18, 25, 26, 27]. These inequalities have been extended to more general networks ([28, 29, 30, 31, 32], see also the review [23]). However, in analogy with the need to define genuine multipartite nonlocality as a means to guarantee global non-classical phenomena, the observation of network nonlocality only guarantees non-classical behavior in some part of the network. It is possible to guarantee a complete absence of classical components via observing genuine multipartite nonlocality, but it is known that accounting for the network structure brings looser requirements for observations [33], which becomes relevant in experimental implementations. Currently, there exist two complementary definitions of genuinely nonlocal behaviors in networks: genuine network nonlocality [34] and full network nonlocality (FNN) [35]. The latter uses minimal assumptions, and has been recently observed in experimental implementations of different network structures [36, 37, 38, 39].
In this work, we study fully network nonlocal correlations under the assumption that physical systems are ultimately governed by quantum mechanics. We derive new Bell-like inequalities, both linear and non-linear, for star and chain networks (see Figs. 1(b) and (c)). The violation of these inequalities certifies that the corresponding statistics could not have been produced had there been at least a number of classical sources in the network. In many cases, we are able to provide a single inequality that guarantees that no source in the network distributes classical physical systems, in contrast with previous works [35, 38] that needed one inequality per source in the network. In other cases, we provide bounds for each number of classical sources in the network. In addition to this, we provide FNN bounds for the inequalities, and demonstrate violations of both types of bounds within quantum mechanics. Finally, we discuss how these results can be exploited to guarantee non-classicality of all the sources of arbitrary networks.
1 Models for non-full correlations
@@
where $\vec{a}=a_{1}\dots a_{n}$ denotes the parties’ outputs, $\vec{x}=x_{1}\dots x_{n}$ their inputs, $\lambda_{j}$ the classical variable distributed by source $j$, and $\bar{\lambda}_{i}$ the set of variables that arrive to party $i$. As an illustration, take the network known as the bilocality scenario, depicted in Fig. 1(a). Network local correlations in this network (i.e., bilocal correlations) admit a description of the form [18, 25]
$P(a,b,c|x,y,z)=\int\text{d}\mu(\lambda_{1})\text{d}\mu(\lambda_{2})P(a|x,\lambda_{1})P(b|y,\lambda_{1},\lambda_{2})P(c|z,\lambda_{2}).$ (2)
@@
Many works have focused on detecting when multipartite correlations do and do not admit realizations in the form of Eq. (1), producing many Bell-like inequalities [18, 25, 28, 29, 30, 31, 32, 23]. However, in generalizing the standard notion of a Bell inequality, inequalities tailored for detecting network nonlocality can be violated even when just one of the sources in the network is non-classical [19]. This motivated the development of new notions beyond network nonlocality whose observation guarantees stronger quantum properties for instance the absence of sources of classical systems [35] or the presence of joint quantum measurements [34] in the network. Concretely, Ref. [35] defines as “not interesting” correlations (i.e., the analogous of Eq. (1) in the setting of network nonlocality) those that can be generated if and only if there is at least one source in the network that distributes classical physical systems. Correlations that cannot be generated in this way are called FNN.
@@
A multipartite probability distribution is $\ell$-level network nonlocal ($\ell$-NN) relative to a network if and only if it cannot be generated when at least $\ell$ sources in the network distribute classical physical systems, and the rest are allowed to distribute physical systems only limited by no-signaling. Otherwise, the distribution is $\ell$-level network local ($\ell$-NL).
@@
$P^{\text{2-NL}}(a_{1},a_{2},a_{3},b|x_{1},x_{2},x_{3})$ $=$
$\int\text{d}\mu(\lambda_{1})\text{d}\mu($ $\lambda_{2})P(a_{1}|x_{1},\lambda_{1})P(a_{2}|x_{2},\lambda_{2})P_{\text{NSI}}(a_{3},b|x_{3},y,\lambda_{1},\lambda_{2}),$ (3)
$P^{\text{1-NL}}(a_{1},a_{2},a_{3},b|x_{1},x_{2},x_{3})$ $=\int\text{d}\mu(\lambda_{1})P(a_{1}|x_{1},\lambda_{1})P_{\text{NSI}}(a_{2},a_{3},b|x_{2},x_{3},y,\lambda_{1}),$ (4)
or of analogues under permutations of the branch parties. $P_{\text{NSI}}$ are distributions only constrained by the principles of no-signaling and the independence of the sources [40]. This means that for the distribution in Eq. (3) we have that $\sum_{a_{3}}P_{\text{NSI}}(a_{3},b|x_{3},y,\lambda_{1},\lambda_{2})=P_{\text{NSI}}(b|y,\lambda_{1},\lambda_{2})$ and $\sum_{b}P_{\text{NSI}}(a_{3},b|x_{3},y,\lambda_{1},\lambda_{2})=P(a_{3}|x_{3})$, and for the case of the distribution in Eq. (4) we require that $\sum_{a_{i}}P_{\text{NSI}}(a_{2},a_{3},b|x_{2},x_{3},y,\lambda_{1})=P_{\text{NSI}}(a_{j\not=i},b|x_{j\not=i},y,\lambda_{1})$ and $\sum_{b}P_{\text{NSI}}(a_{2},a_{3},b|x_{2},x_{3},y,\lambda_{1})=P(a_{2}|x_{2})P(a_{3}|x_{3})$.
Definition 1 allows to make theory-independent statements on the presence of classical sources in networks [38]. However, inserting the additional assumption that the non-classical sources in the network distribute systems that obey quantum theory is a natural one that, as we will show, leads to milder requirements on the experimentally relevant parameters (such as the visibility under white noise) needed to observe nonlocality. Thus, we also study the hierarchy in Definition 1 under this assumption. This leads to an analogous hierarchy, and to the definition of fully quantum network nonlocality (FQNN):
@@
A multipartite probability distribution is $\ell$-level quantum network nonlocal ($\ell$-QNN) relative to a network if and only if it cannot be generated when at least $\ell$ sources in the network distribute classical physical systems, and the rest are allowed to distribute arbitrary multipartite quantum systems. Otherwise, the distribution is $\ell$-level quantum network local ($\ell$-QNL). If $\ell=1$, we say that the correlations are fully quantum network nonlocal.
Going back to the example of the star network with three branch parties, the corresponding distributions $P^{\text{2-QNL}}$ and $P^{\text{1-QNL}}$ have the same form as Eqs. (3), (4), where now the distributions $P_{\text{NSI}}$ are instead distributions that follow the corresponding Born’s rule, i.e. $\text{Tr}(\Pi_{a_{3}}^{x_{3}}\otimes\Pi_{b}^{y,\lambda_{1},\lambda_{2}}\cdot\rho_{A_{3}B})$ for Eq. (3) and $\text{Tr}(\Pi_{a_{2}}^{x_{2}}\otimes\Pi_{a_{3}}^{x_{3}}\otimes\Pi_{b}^{y,\lambda_{1}}\cdot\rho_{A_{2}B}\otimes\rho_{A_{3}B})$ for Eq. (4).
In the remainder of this manuscript we analytically prove inequalities, both linear and non-linear, satisfied by $\ell$-NL and $\ell$-QNL correlations in large families of networks and for several values of $\ell$, including complete hierarchies (i.e., $\ell\in[1,\dots,m]$) for some scenarios. We must note that, in general, obtaining analytical results for $\ell$-NL is hard. This is due to the lack of a formal description of all generalized states and effects that lead to non-signaling distributions. For a more systematic analysis than the one done in this work one could use inflation methods (see, e.g., [35, 38] for examples of its use in the derivation of witnesses of FNN), to create tractable relaxations of the sets of $\ell$-NL distributions.
Figure 2: The $n$-partite chain network. Correlations that do not admit a realization neither in the form of (a) nor with any other positioning of the classical variable are FQNN, and those that do not admit a realization neither in the form of (b) nor with any other positioning of the classical variable are FNN. There exist correlations of both types that can be produced in the setup depicted in (c).
2 Nonlinear inequalities for chain networks
The first set of networks that we study is that of chain networks, depicted in Fig. 1(b). Chain networks are comprised of sources of bipartite states with the parties arranged in the form of a chain, such that with the exception of the two, extremal parties, all the remaining ones receive systems from two different sources. We will initially analyze the simplest case of the tripartite chain network, and extend our results to longer chains in a second step.
2.1 The tripartite chain
@@
where $I$ and $J$ are defined by $I=\frac{1}{4}(A_{0}B^{0}C_{0}+A_{1}B^{0}C_{0}+A_{0}B^{0}C_{1}+A_{1}B^{0}C_{1})$ and $J=\frac{1}{4}(A_{0}B^{1}C_{0}-A_{1}B^{1}C_{0}-A_{0}B^{1}C_{1}+A_{1}B^{1}C_{1})$, $A_{i}$ and $C_{k}$ are the (dichotomic) measurement operators of the extreme parties, $B$ denotes one measurement with two-bit outcomes, depicted $b^{0}b^{1}$, and the correlator $A_{x}B^{y}C_{z}$ is given by $\sum_{a,b^{0}b^{1},c}(-1)^{a+b^{y}+c}P(a,b^{0}b^{1},c|x,z)$.
@@
$\sqrt{|I|}+\sqrt{|J|}=\,$ $\frac{1}{2}\sqrt{|A_{0}B^{0}C_{0}+A_{1}B^{0}C_{0}+A_{0}B^{0}C_{1}+A_{1}B^{0}C_{1}|}$
$+\frac{1}{2}\sqrt{|A_{0}B^{1}C_{0}-A_{1}B^{1}C_{0}-A_{0}B^{1}C_{1}+A_{1}B^{1}C_{1}|}$
@@
Now, assume that $A_{0}+A_{1}\geq 0$ and define $x=A_{0}+A_{1}-1$, so that $x\in[-1,1]$. Note that $|A_{0}-A_{1}|+|A_{0}+A_{1}|\leq 2\max\{|A_{0}|,|A_{1}|\}\leq 2$. Therefore, we have that $|A_{0}-A_{1}|\leq 2-|A_{0}+A_{1}|=1-x$. Using this, Eq. (6) reads
$\sqrt{|I|}+\sqrt{|J|}\leq\,$ $\frac{1}{2}\left(\sqrt{(1+x)|B^{0}C_{0}+B^{0}C_{1}|}+\sqrt{(1-x)|B^{1}C_{0}-B^{1}C_{1}|}\right)$
=\,= $\frac{1}{\sqrt{2}}\left(\cos\theta\sqrt{|B^{0}C_{0}+B^{0}C_{1}|}+\sin\theta\sqrt{|B^{1}C_{0}-B^{1}C_{1}|}\right)$ (7)
[…]
where in Eq. (7) we have rewritten $1+x:=2\cos^{2}\theta$ and $1-x:=2\sin^{2}\theta$, and in Eq. (8) we have used that $a\cos\theta+b\sin\theta\leq\sqrt{a^{2}+b^{2}}$. If, instead, $A_{0}+A_{1}<0$, one can define $x=A_{0}+A_{1}+1$ and the same result follows.
Next, note for a given value $y\in\{0,1\}$, the object $B^{y}$ only considers the $y$-th bit of Bob’s output. One can thus view the correlators $B^{y}C_{z}$ as generated by the following two-input, two-output distribution:
$P(b,c|y,z)=P(b=b^{y},c|z)=\sum_{b^{0},b^{1}}\delta_{b,b^{y}}P(b^{0}b^{1},c|z),$
in which case it follows that
@@
Therefore, the bounds on $\sqrt{|I|}+\sqrt{|J|}$ for FNN and FQNN are connected to the maximum value of the CHSH inequality [41] that Bob and Charlie can achieve. When Bob and Charlie share a classical physical system, $\text{CHSH}(B,C)\leq 2$ and thus any value $\sqrt{|I|}+\sqrt{|J|}>\sqrt{2}/\sqrt{2}=1$ implies standard network nonlocality, as derived in Refs. [18, 25]. If they share an arbitrary system only limited by the no-signaling principle, then $\text{CHSH}(B,C)\leq 4$, and any value $\sqrt{|I|}+\sqrt{|J|}>\sqrt{4}/\sqrt{2}=\sqrt{2}$ implies FNN, as derived in Ref. [35]. Finally, if the parties share a quantum system, then $\text{CHSH}(B,C)\leq 2\sqrt{2}$ and any value $\sqrt{|I|}+\sqrt{|J|}>\sqrt{2\sqrt{2}}/\sqrt{2}=2^{1/4}$ implies FQNN.
The same argumentation can be done for the case where the classical source is shared between Bob and Charlie, arriving to the same inequality, $\sqrt{|I|}+\sqrt{|J|}\leq 2^{1/4}$. Therefore, this is a single inequality whose violation guarantees that none of the sources in the scenario is classical. This is in contrast to earlier works on detection of FNN [35, 38], which needed to violate as many inequalities as sources in the network in order to guarantee that all of them were non-classical. Moreover, both bounds are tight: the bound for FQNN is saturated by the strategy in Ref. [42] and the bound for FNN is saturated by the strategy in Ref. [35]. In fact, the bound for FNN coincides with the algebraic maximum of $\sqrt{|I|}+\sqrt{|J|}$, computed in Ref. [35]. This means that $\sqrt{|I|}+\sqrt{|J|}$ cannot be used for detecting FNN, but it can be used for detecting FQNN. Indeed, take the sources to distribute pure states, $\rho_{i}=|\phi_{i}\rangle\langle\phi_{i}|$, given by
@@
where $\theta_{i}\in(0,\frac{\pi}{2})$. Let also $A_{x}=C_{z}=\cos\vartheta\sigma_{3}+(-1)^{x}\sin\vartheta\sigma_{1}$ and $B^{b^{0}b^{1}}\in\{\Phi_{0},\Phi_{1},\Psi_{0},\Psi_{1}\}$, where $\sigma_{1}$ and $\sigma_{3}$ are the Pauli matrices, and $\rho=|\rho\rangle\langle\rho|$ with $|\Phi_{i}\rangle=(|00\rangle+(-1)^{i}|11\rangle)/\sqrt{2}$ and $|\Psi_{i}\rangle=(|01\rangle+(-1)^{i}|10\rangle)/\sqrt{2}$ being the Bell states. A straightforward evaluation leads to

### v4

# Hierarchical certification of nonclassical network correlations

@@ Abstract
With the increased availability of quantum technological devices, it becomes more important to have tools to guarantee their correct nonclassical behavior. This is especially important for quantum networks, which constitute the platforms where multipartite cryptographic protocols will be implemented, and where guarantees of nonclassicality translate into security proofs. We derive linear and nonlinear Bell-like inequalities for networks, whose violation certifies the absence of a minimum number of classical sources in them. We do so, first, without assuming that nature is ultimately governed by quantum mechanics, providing a hierarchy interpolating between network nonlocality and full network nonlocality. Second we insert this assumption, which leads to results more amenable to certification in experiments.
I Introduction
@@
Figure 1: The main networks discussed in this work. (a) The network underlying simple entanglement swapping experiments. Correlations generated in the network are considered NN if they cannot be generated as in the top-left setup, FNN if they cannot be generated as in the bottom-left setup, and FQNN if they cannot be generated as in the top-right setup. In the appendices we show that realizations in the bottom-right setup produce all these types of correlations. (b) The generalization of the previous scenario to chains, and (c) its generalization to stars.
Nonlinear Bell inequalities have been used to study network nonlocality (NN), especially in the tripartite network underlying entanglement swapping depicted in Fig. 1(a) [18, 25, 26, 27]. These inequalities have been extended to more general networks with more parties and sources of physical systems ([28, 29, 30, 31, 32], see also the review [23]). However, in analogy with the need to define genuine multipartite nonlocality as a means to guarantee global nonclassical phenomena, the observation of network nonlocality only guarantees nonclassical behavior in some part of the network. It is possible to guarantee a complete absence of classical components via observing genuine multipartite nonlocality, but it is known that accounting for the network structure allows for milder requirements in experimentally relevant parameters, such as detection efficiency [33]. Currently, there exist two complementary definitions of genuinely nonlocal behaviors in networks: genuine network nonlocality [34] and full network nonlocality (FNN) [35]. The latter uses minimal assumptions, and has been recently observed in experimental implementations of different network structures [36, 37, 38, 39].
In this work, we study fully network nonlocal correlations under the assumption that physical systems are ultimately governed by quantum mechanics. We derive new Bell-like inequalities, both linear and nonlinear, for star and chain networks (see Figs. 1(b) and (c)). The violation of these inequalities certifies that the corresponding statistics could not have been produced had there been at least a number of classical sources in the network. In many cases, we are able to provide a single inequality that guarantees that no source in the network distributes classical physical systems, in contrast with previous works [35, 38] that needed one inequality per source in the network. In other cases, we provide bounds for each number of classical sources in the network. In addition to this, we provide FNN bounds for the inequalities, and demonstrate violations of both types of bounds within quantum mechanics. Finally, we discuss how these results can be exploited to guarantee nonclassicality of the sources of arbitrary networks.
II Models for non-full correlations
@@
where $\vec{a}=a_{1}\dots a_{n}$ denotes the parties’ outputs, $\vec{x}=x_{1}\dots x_{n}$ their inputs, $\lambda_{j}$ the classical variable distributed by source $j$, and $\bar{\lambda}_{i}$ the set of variables that arrive to party $i$, i.e., the set of classical physical systems that arrive to party $i$ from the sources that connect to it. As an illustration, take the network known as the bilocality scenario, depicted in Fig. 1(a). Network-local correlations in this network (i.e., bilocal correlations) admit a description of the form [18, 25]
$P(a,b,c|x,y,z)=\int$ $\text{d}\mu(\lambda_{1})\text{d}\mu(\lambda_{2})$
$P(a|x,\lambda_{1})P(b|y,\lambda_{1},\lambda_{2})P(c|z,\lambda_{2}).$
@@
Many works have focused on detecting when multipartite correlations do and do not admit realizations in the form of Eq. (1), producing many Bell-like inequalities [18, 25, 28, 29, 30, 31, 32, 23]. However, in generalizing the standard notion of a Bell inequality, inequalities tailored for detecting network nonlocality can be violated even when just one of the sources in the network is nonclassical [19]. This motivated the development of new notions beyond network nonlocality whose observation guarantees stronger quantum properties, for instance the absence of sources of classical systems [35] or the presence of joint quantum measurements [34] in the network. Concretely, Ref. [35] defines as “not interesting” correlations (i.e., the analogous of Eq. (1) in the setting of network nonlocality) those that can be generated if and only if there is at least one source in the network that distributes classical physical systems. Correlations that cannot be generated in this way are called FNN.
@@
A multipartite probability distribution is $\ell$-level network nonlocal ($\ell$-NN) relative to a network if it cannot be generated when at least $\ell$ sources in the network distribute classical physical systems, and the rest are allowed to distribute physical systems only limited by no-signaling. Otherwise, the distribution is $\ell$-level network local ($\ell$-NL).
@@
$P^{\text{2-NL}}($ $a_{1},a_{2},a_{3},b|x_{1},x_{2},x_{3},y)=\int\text{d}\mu(\lambda_{1})\text{d}\mu(\lambda_{2})$
$P(a_{1}|x_{1},\lambda_{1})P(a_{2}|x_{2},\lambda_{2})P_{\text{NSI}}(a_{3},b|x_{3},y,\lambda_{1},\lambda_{2}),$ (2)
$P^{\text{1-NL}}($ $a_{1},a_{2},a_{3},b|x_{1},x_{2},x_{3},y)=$
[…]
or of analogs under permutations of the branch parties. $P_{\text{NSI}}$ are distributions only constrained by the principles of no-signaling and the independence of the sources [40]. This means that for the distribution in Eq. (2) we have that $\sum_{a_{3}}P_{\text{NSI}}(a_{3},b|x_{3},y,\lambda_{1},\lambda_{2})=P_{\text{NSI}}(b|y,\lambda_{1},\lambda_{2})$ and $\sum_{b}P_{\text{NSI}}(a_{3},b|x_{3},y,\lambda_{1},\lambda_{2})=P(a_{3}|x_{3})$, and for the case of the distribution in Eq. (3) we require that $\sum_{a_{i}}P_{\text{NSI}}(a_{2},a_{3},b|x_{2},x_{3},y,\lambda_{1})=P_{\text{NSI}}(a_{j\not=i},b|x_{j\not=i},y,\lambda_{1})$ and $\sum_{b}P_{\text{NSI}}(a_{2},a_{3},b|x_{2},x_{3},y,\lambda_{1})=P(a_{2}|x_{2})P(a_{3}|x_{3})$.
Definition 1 allows to make theory-independent statements on the presence of classical sources in networks [38]. However, inserting the additional assumption that the nonclassical sources in the network distribute systems that obey quantum theory is a natural one that, as we will show, leads to milder requirements on the experimentally relevant parameters (such as the visibility under white noise) needed to observe nonlocality. Thus, we also study the hierarchy in Definition 1 under this assumption. This leads to an analogous hierarchy, and to the definition of fully quantum network nonlocality (FQNN):
@@
A multipartite probability distribution is $\ell$-level quantum network nonlocal ($\ell$-QNN) relative to a network if it cannot be generated when at least $\ell$ sources in the network distribute classical physical systems, and the rest are allowed to distribute arbitrary multipartite quantum systems. Otherwise, the distribution is $\ell$-level quantum network local ($\ell$-QNL). We say that a correlation is fully quantum network nonlocal if it is 1-QNN.
Figure 2: The $n$-partite chain network. Correlations that do not admit a realization neither in the form of panel (a) nor with any other positioning of the classical variable are FQNN, and those that do not admit a realization neither in the form of panel (b) nor with any other positioning of the classical variable are FNN. There exist correlations of both types that can be produced in the setup depicted in panel (c).
Going back to the example of the star network with three branch parties, the corresponding distributions $P^{\text{2-QNL}}$ and $P^{\text{1-QNL}}$ have the same form as Eqs. (2) and (3), where now the distributions $P_{\text{NSI}}$ are instead distributions that follow the corresponding Born’s rule, i.e. $\text{Tr}(\Pi_{a_{3}}^{x_{3}}\otimes\Pi_{b}^{y,\lambda_{1},\lambda_{2}}\cdot\rho_{A_{3}B})$ for Eq. (2) and $\text{Tr}(\Pi_{a_{2}}^{x_{2}}\otimes\Pi_{a_{3}}^{x_{3}}\otimes\Pi_{b}^{y,\lambda_{1}}\cdot\rho_{A_{2}B}\otimes\rho_{A_{3}B})$ for Eq. (3).
In the remainder of this manuscript we analytically prove inequalities, both linear and nonlinear, satisfied by $\ell$-NL and $\ell$-QNL correlations in large families of networks and for several values of $\ell$, including complete hierarchies (i.e., $\ell\in[1,\dots,m]$) for some scenarios. We must note that, in general, obtaining analytical results for $\ell$-NL is hard. This is due to the lack of a formal description of all generalized states and effects that lead to nonsignaling distributions. For a more systematic analysis than the one done in this work one could use inflation methods (see, e.g., [35, 38, 59] for examples of its use in the derivation of witnesses of FNN), to create tractable relaxations of the sets of $\ell$-NL distributions.
III Nonlinear inequalities for chain networks
The first family of networks that we study is that of chain networks, depicted in Figs. 1(b) and 2. Chain networks are comprised of sources of bipartite states with the parties arranged in the form of a chain, such that with the exception of the two, extremal parties, all the remaining ones receive systems from two different sources. We will initially analyze the simplest case of the tripartite chain network, and extend our results to longer chains in a second step.
III.1 The tripartite chain
@@
where $I$ and $J$ are defined by $I=\frac{1}{4}\left\langle A_{0}B^{0}C_{0}+A_{1}B^{0}C_{0}+A_{0}B^{0}C_{1}+A_{1}B^{0}C_{1}\right\rangle$ and $J=\frac{1}{4}\left\langle A_{0}B^{1}C_{0}-A_{1}B^{1}C_{0}-A_{0}B^{1}C_{1}+A_{1}B^{1}C_{1}\right\rangle$, $A_{i}$ and $C_{k}$ are the (dichotomic) measurement operators of the extreme parties, $B$ denotes one measurement with two-bit outcomes, depicted $b^{0}b^{1}$, and the correlator $\left\langle A_{x}B^{y}C_{z}\right\rangle$ is given by $\sum_{a,b^{0}b^{1},c}(-1)^{a+b^{y}+c}P(a,b^{0}b^{1},c|x,z)$.
@@
$\sqrt{|I|}+\sqrt{|J|}=\,$ $\frac{1}{2}\sqrt{|\left\langle A_{0}B^{0}C_{0}+A_{1}B^{0}C_{0}+A_{0}B^{0}C_{1}+A_{1}B^{0}C_{1}\right\rangle|}+\frac{1}{2}\sqrt{|\left\langle A_{0}B^{1}C_{0}-A_{1}B^{1}C_{0}-A_{0}B^{1}C_{1}+A_{1}B^{1}C_{1}\right\rangle|}$
@@
Now, assume that $\left\langle A_{0}+A_{1}\right\rangle\geq 0$ and define $x=\left\langle A_{0}+A_{1}\right\rangle-1$, so that $x\in[-1,1]$. Note that $|\left\langle A_{0}-A_{1}\right\rangle|+|\left\langle A_{0}+A_{1}\right\rangle|\leq 2\max\{|\left\langle A_{0}\right\rangle|,|\left\langle A_{1}\right\rangle|\}\leq 2$. Therefore, we have that $|\left\langle A_{0}-A_{1}\right\rangle|\leq 2-|\left\langle A_{0}+A_{1}\right\rangle|=1-x$. Using this, Eq. (5) reads
$\sqrt{|I|}+\sqrt{|J|}\leq\,$ $\frac{1}{2}\left(\sqrt{(1+x)|\left\langle B^{0}C_{0}+B^{0}C_{1}\right\rangle|}\right.$
$\left.\quad+\sqrt{(1-x)|\left\langle B^{1}C_{0}-B^{1}C_{1}\right\rangle|}\right)$
=\,= $\frac{1}{\sqrt{2}}\left(\cos\theta\sqrt{|\left\langle B^{0}C_{0}+B^{0}C_{1}\right\rangle|}\right.$
$\left.\qquad+\sin\theta\sqrt{|\left\langle B^{1}C_{0}-B^{1}C_{1}\right\rangle|}\right)$ (6)
$\leq\,\frac{1}{\sqrt{2}}\sqrt{|\left\langle B^{0}C_{0}+B^{0}C_{1}\right\rangle+|\left\langle B^{1}C_{0}-B^{1}C_{1}\right\rangle|},$ (7)
where in Eq. (6) we have rewritten $1+x:=2\cos^{2}\theta$ and $1-x:=2\sin^{2}\theta$, and in Eq. (7) we have used that $a\cos\theta+b\sin\theta\leq\sqrt{a^{2}+b^{2}}$. If, instead, $\left\langle A_{0}+A_{1}\right\rangle<0$, one can define $x=\left\langle A_{0}+A_{1}\right\rangle+1$ and the same result follows.
Next, note that the object $B^{y}$ only considers the $y$-th bit of Bob’s output. One can thus view the correlators $\left\langle B^{y}C_{z}\right\rangle$ as generated by the following simulated, two-input, two-output distribution:
$P_{\text{sim}}(b,c|y,z)=P(b=b^{y},c|z)=\sum_{b^{0},b^{1}}\delta_{b,b^{y}}P(b^{0}b^{1},c|z),$
which is obtained by Bob outputting only the $y$-th bit upon choosing measurement $y$. Moreover, it is easy to see that $|\left\langle B^{0}C_{0}+B^{0}C_{1}\right\rangle|+|\left\langle B^{1}C_{0}-B^{1}C_{1}\right\rangle|$ is the largest of $|\left\langle B^{0}C_{0}+B^{0}C_{1}+(B^{1}C_{0}-B^{1}C_{1})\right\rangle|$ and $|\left\langle B^{0}C_{0}+B^{0}C_{1}-(B^{1}C_{0}-B^{1}C_{1})\right\rangle|$. These are two variants of the CHSH operator under relabeling of Charlie’s input. Thus, it follows that
@@
Therefore, the bounds on $\sqrt{|I|}+\sqrt{|J|}$ for FNN and FQNN are connected to the maximum value of the CHSH inequality [41] that Bob and Charlie can achieve. When Bob and Charlie share a classical physical system, $\text{CHSH}(B,C)\leq 2$ and thus any value $\sqrt{|I|}+\sqrt{|J|}>\sqrt{2}/\sqrt{2}=1$ implies standard network nonlocality, as derived in Refs. [18, 25]. If they share an arbitrary system only limited by the no-signaling principle, then $\text{CHSH}(B,C)\leq 4$, and any value $\sqrt{|I|}+\sqrt{|J|}>\sqrt{4}/\sqrt{2}=\sqrt{2}$ implies FNN, as derived in Ref. [35]. Finally, if the parties share a quantum system, then $\text{CHSH}(B,C)\leq 2\sqrt{2}$ and any value $\sqrt{|I|}+\sqrt{|J|}>\sqrt{2\sqrt{2}}/\sqrt{2}=2^{1/4}$ implies FQNN. This is a proof that the bound first provided in Ref. [42] is tight.
The same argumentation can be done for the case where the classical source is shared between Bob and Charlie, arriving to the same inequality, $\sqrt{|I|}+\sqrt{|J|}\leq 2^{1/4}$. Therefore, this is a single inequality whose violation guarantees that none of the sources in the scenario is classical. This is in contrast to earlier works on detection of FNN [35, 38], which needed to violate as many inequalities as sources in the network in order to guarantee that all of them were nonclassical. Moreover, both bounds are tight: the bound for FQNN is saturated by the strategy in Ref. [42] and the bound for FNN is saturated by the strategy in Ref. [35]. In fact, the bound for FNN coincides with the algebraic maximum of $\sqrt{|I|}+\sqrt{|J|}$, computed in Ref. [35]. This means that $\sqrt{|I|}+\sqrt{|J|}$ cannot be used for detecting FNN, but it can be used for detecting FQNN. Indeed, take the sources to distribute pure states, $\rho_{i}=|\phi_{i}\rangle\langle\phi_{i}|$, given by
@@
where $\theta_{i}\in(0,\frac{\pi}{2})$. Let also $A_{x}=\cos\vartheta\sigma_{3}+(-1)^{x}\sin\vartheta\sigma_{1}$ and likewise for $C_{z}$, and $B^{b^{0}b^{1}}\in\{\Phi_{0},\Phi_{1},\Psi_{0},\Psi_{1}\}$, where $\sigma_{1}$ and $\sigma_{3}$ are the Pauli matrices, and $\rho=|\rho\rangle\langle\rho|$ with $|\Phi_{i}\rangle=(|00\rangle+(-1)^{i}|11\rangle)/\sqrt{2}$ and $|\Psi_{i}\rangle=(|01\rangle+(-1)^{i}|10\rangle)/\sqrt{2}$ being the Bell states. A straightforward evaluation leads to